# Learning to Track Any Points from Human Motion

### v1

@@ 3.1 Point Tracking with Human Mesh Models
Conversely, the vertex $\mathbf{v}_{p,t,j}$ is considered occluded ($v_{p,t,j}=0$) if the closest intersection point along the ray (the one with the minimum $s_{\text{int},k}$) is with a triangle belonging to any human mesh surface, and that intersection is strictly closer to the camera than $\mathbf{v}_{p,t,j}$ itself, i.e., $\min_{k,\forall p^{\prime}}(s_{\text{int},k})<1$. This ray-triangle intersection testing is efficiently implemented using the Möller-Trumbore algorithm [43], which can be significantly accelerated on modern GPUs. Note that this ray casting method accounts for occlusions by the modeled human bodies (both self and inter-person) and does not detect occlusions by other arbitrary scene objects. Limitations regarding occlusions by non-modeled scene elements are addressed in the subsequent filtering stage.
@@
3.2 Filtering Erroneous Tracks with Optical Flow
While HMR methods [5, 17, 20] have become increasingly robust to challenges such as motion blur, rapid movements, and varying lighting conditions due to their strong human priors, they may still produce occasional errors or inconsistencies in mesh predictions under extreme conditions. Furthermore, our ray-casting-based visibility prediction is limited to occlusions caused by other modeled humans and does not account for occlusions from general scene objects. In addition, the parametric nature of the SMPL model makes it less effective at capturing the motion of highly deformable or loosely attached clothing and accessories.
To mitigate the impact of these potential inaccuracies on our pseudo-labels and thus improve the quality of training data, we introduce an optical flow-based filtering stage, illustrated in Figure 3. Optical flow [58, 24, 65] is particularly well-suited for this task due to its established accuracy in estimating dense motion between adjacent frames. Over very short temporal windows, these flow-based predictions can provide strong local motion cues to validate or identify discrepancies in the HMR-derived trajectories, especially when dealing with complex non-rigid deformations or potential tracking drift [11, 44].
@@ 3.2 Filtering Erroneous Tracks with Optical Flow
Identify confident optical flow. For each pair of consecutive frames $(I_{t},I_{t+1})$, we compute both the forward optical flow map from $I_{t}$ to $I_{t+1}$, denoted $\mathbf{F}_{t\to t+1}$, and the backward optical flow map $\mathbf{F}_{t+1\to t}$. We use $\mathbf{f}(\mathbf{F},\mathbf{x})$ to denote the process of sampling the flow vector from map $\mathbf{F}$ at sub-pixel location $\mathbf{x}$ using bilinear interpolation. To assess the reliability of the optical flow estimation itself at a given point $\mathbf{x}_{p,t,j}$, we apply a forward-backward consistency check [42]. Let $c_{p,t,j}$ denote the binary indicator of optical flow reliability at point $\mathbf{x}_{p,t,j}$ in frame $I_{t}$. A pixel $\mathbf{x}_{p,t,j}$ is considered to have reliable flow ($c_{p,t,j}=1$) if the $L_{2}$ distance between the point warped to frame $t+1$ and then back to frame $t$ is below a predefined threshold $\delta_{\text{cons}}$:
$c_{p,t,j}=\mathbbm{1}\left[\|\mathbf{f}(\mathbf{F}_{t\to t+1},\mathbf{x}_{p,t,j})+\mathbf{f}(\mathbf{F}_{t+1\to t},\mathbf{x}_{p,t,j}+\mathbf{f}(\mathbf{F}_{t\to t+1},\mathbf{x}_{p,t,j}))\|_{2}<\delta_{\text{cons}}\right].$ (2)
Find erroneous trajectories. For each point $\mathbf{x}_{p,t,j}$ within a pseudo-labeled trajectory $\mathcal{X}_{p,j}$ (where $t$ and $t+1$ are both in $\mathcal{T}_{p,j}$), we examine the predicted motion to the next frame, $\mathbf{x}_{p,t+1,j}$. The displacement vector derived from the HMR pseudo-label is $\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}=\mathbf{x}_{p,t+1,j}-\mathbf{x}_{p,t,j}$. We compare this to the optical flow displacement at the location $\mathbf{x}_{p,t,j}$, obtained by sampling the forward optical flow map: $\Delta\mathbf{x}_{p,t,j}^{\text{OF}}=\mathbf{f}(\mathbf{F}_{t\to t+1},\mathbf{x}_{p,t,j})$.
To compare these two displacement vectors, $\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}$ and $\Delta\mathbf{x}_{p,t,j}^{\text{OF}}$, robustly, especially in cases of large motion, we normalize them. Let $L_{\text{short}}=\min(\|\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}\|_{2},\|\Delta\mathbf{x}_{p,t,j}^{\text{OF}}\|_{2})+\epsilon_{\text{norm}}$, where $\epsilon_{\text{norm}}$ is a small positive constant added to prevent division by zero. The normalized vectors are:
$\Delta\widehat{\mathbf{x}}_{p,t,j}^{\text{HMR}}=\frac{\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}}{L_{\text{short}}},\quad\Delta\widehat{\mathbf{x}}_{p,t,j}^{\text{OF}}=\frac{\Delta\mathbf{x}_{p,t,j}^{\text{OF}}}{L_{\text{short}}}.$ (3)
@@
$e_{p,t,j}=\mathbbm{1}\left[\|\Delta\widehat{\mathbf{x}}_{p,t,j}^{\text{HMR}}-\Delta\widehat{\mathbf{x}}_{p,t,j}^{\text{OF}}\|_{2}>\tau_{\text{dist}}\right].$ (4)

### v2

@@ 3.1 Point tracking with human mesh models
Conversely, the vertex $\mathbf{v}_{p,t,j}$ is considered occluded ($v_{p,t,j}=0$) if the closest intersection point along the ray (the one with the minimum $s_{\text{int},k}$) is with a triangle belonging to any human mesh surface, and that intersection is strictly closer to the camera than $\mathbf{v}_{p,t,j}$ itself, i.e., $\min_{k,\forall p^{\prime}}(s_{\text{int},k})<1$. This ray-triangle intersection testing is efficiently implemented using the Möller-Trumbore algorithm [moller2005fast], which can be significantly accelerated on modern GPUs. Note that this ray casting method accounts for occlusions by the modeled human bodies (both self and inter-person) and does not detect occlusions by other arbitrary scene objects. Limitations regarding occlusions by non-modeled scene elements are addressed in the subsequent filtering stage.
3.2 Supervised training with pseudo-labels
Vertex projection and ray casting for visibility prediction yield a collection of refined 2D point trajectories $\mathcal{X}_{p,j}$ with associated with visibility labels $v_{p,j}$. These trajectories serve as additional supervision for existing point tracking models [cho2024local, karaev2024cotracker3], and we simply apply each model’s original training loss. To further improve the quality of our pseudo-labeled human tracks, we introduce an additional optical flow–based filtering during the training that identifies and removes unreliable trajectory segments.
Optical-flow based trajectory filtering. Although modern HMR methods are generally robust, they still fail under extreme poses or occlusions. Since optical flow [teed2020raft] provides accurate short-term motion cues, we compare the HMR-predicted motion of each tracked point with the corresponding flow-derived displacement.
For each trajectory, we identify transitions whose motion disagrees with the optical flow beyond a normalized threshold, where a forward–backward consistency check is used to ensure that the sampled flow is locally reliable. We then compute the fraction of erroneous transitions, considering only frames where both the HMR visibility and optical-flow reliability conditions are met. If this ratio exceeds a preset limit, the entire trajectory is discarded; otherwise, only the inconsistent transitions are removed. This process ensures that the final training set contains fewer but much more accurate pseudo-labels, which is preferable for learning a robust point tracker. The full details can be found in Sec. C of the supplementary material.
Table 1: Quantitative results on TAP-Vid benchmark [doersch2022tap] and RoboTAP [vecerik2024robotap]. LocoTrack-B, trained with our approach, shows a significant performance improvement across all metrics and datasets. Furthermore, in terms of position accuracy ($\delta_{avg}^{x}$), our model achieves the best performance on every dataset. We use a training dataset that is $11\times$ smaller than CoTracker3 [karaev2024cotracker3] in terms of the number of videos, and $1{,}000\times$ smaller than the dataset used in BootsTAPIR [doersch2024bootstap] in terms of the number of training frames. Kub refers to the Kubric dataset [greff2022kubric], while Kub64 refers to the dataset rendered in Karaev et al. [karaev2024cotracker3]. Both Kub and Kub64 are synthetic datasets.
Method Training DAVIS First DAVIS Strided Kinetics First RoboTAP First [vecerik2024robotap]
Dataset AJ $\delta_{avg}^{x}$ OA AJ $\delta_{avg}^{x}$ OA AJ $\delta_{avg}^{x}$ OA AJ $\delta_{avg}^{x}$ OA
Models evaluated at $256\times 256$ resolution
OmniMotion [wang2023tracking] - - - - $51.7$ $67.5$ $85.3$ - - - - - -
Dino-Tracker [tumanyan2024dino] - - - - $62.3$ $78.2$ $87.5$ - - - - - -
TAPNet [doersch2022tap] Kub $33.0$ $48.6$ $78.8$ $38.4$ $53.1$ $82.3$ $38.5$ $54.4$ $80.6$ - - -
TAPIR [doersch2023tapir] Kub $58.5$ $70.0$ $86.5$ $61.3$ $73.6$ $88.8$ $49.6$ $64.2$ $85.0$ $59.6$ $73.4$ $87.0$
Online TAPIR [vecerik2024robotap] Kub $56.2$ $70.0$ $86.5$ - - - $51.5$ - - - - -
TAPTR [li2024taptr] Kub $63.0$ $76.1$ $\underline{91.1}$ $66.3$ $79.2$ $\underline{91.0}$ $49.0$ $64.4$ $85.2$ $60.1$ $75.3$ $86.9$
TAPTRv2 [li2024taptrv2] Kub 63.5 $75.9$ 91.4 $66.4$ $78.8$ 91.3 $49.7$ $64.2$ $85.7$ - - -
TAPTRv3 [qu2024taptrv3] Kub $63.2$ $\underline{76.7}$ $91.0$ - - - $\underline{54.5}$ $\underline{67.5}$ 88.2 - - -
BootsTAPIR [doersch2024bootstap] Kub+15M $61.4$ $74.0$ $88.4$ $66.2$ $78.5$ $90.7$ 54.6 68.4 $\underline{86.5}$ 64.9 80.1 $86.3$
LocoTrack [cho2024local] Kub $63.0$ $75.3$ $87.2$ $\underline{67.8}$ $\underline{79.6}$ $89.9$ $52.9$ $66.8$ $85.3$ $62.3$ $76.2$ $\underline{87.1}$
Models evaluated at $384\times 512$ resolution
PIPs [harley2022particle] FT [mayer2016large] $42.2$ $64.8$ $77.7$ $52.4$ $70.0$ $83.6$ - - - - - -
CoTracker2 [karaev2024cotracker] Kub $62.2$ $75.7$ $89.3$ $65.9$ $79.4$ $\underline{89.9}$ $48.8$ $64.5$ $85.8$ - - -
Track-On [aydemir2025track] Kub $\underline{65.0}$ $\underline{78.0}$ $\underline{90.8}$ - - - $53.9$ $67.3$ $\underline{87.8}$ - - -
CoTracker3 (online) [karaev2024cotracker3] Kub64+15K $64.4$ $76.9$ 91.2 - - - $54.7$ $67.8$ $87.4$ - - -
CoTracker3 (offline) [karaev2024cotracker3] Kub64+15K $63.8$ $76.3$ $90.2$ - - - 55.8 $\underline{68.5}$ $\mathbf{88.3}$ - - -
LocoTrack [cho2024local] Kub $64.8$ $77.4$ $86.2$ $\underline{69.4}$ $\underline{81.3}$ $88.6$ $52.3$ $66.4$ $82.1$ - - -
Anthro-LocoTrack (Ours) Kub+1.4K 65.9 78.9 $87.3$ $\mathbf{71.1}$ 82.9 90.3 $\underline{54.8}$ 68.6 $85.3$ - - -
4 Experiments
Figure 3: Qualitative results on the videos from DAVIS dataset [pont20172017]. We present comparisons with CoTracker3 [karaev2024cotracker3] and BootsTAPIR [doersch2024bootstap]. Anthro-LocoTrack (Ours) consistently demonstrates strong performance on highly deformable objects and severe occlusions.
4.1 Experimental setup
Dataset construction.
We use the pre-trained TokenHMR [dwivedi2024tokenhmr] model for human mesh recovery and SEA-RAFT [wang2024sea] for obtaining optical flow. To construct our training dataset, we use 1.4K videos from the Let’s Dance dataset [castro2018let], which includes a diverse range of dance performances, from solo acts to multi-person scenes. Unlike BootsTAP [doersch2024bootstap] or CoTracker3 [doersch2024bootstap], our training videos are entirely non-proprietary. This makes the training and annotation process fully reproducible, which we believe will make a valuable contribution to the point tracking community.
Training details. We fine-tune the pre-trained LocoTrack [cho2024local] base model using AnthroTAP-annotated Let’s Dance dataset, which called Anthro-LD. During fine-tuning, we randomly sample data from our dataset and Kubric Panning MOVi-E [doersch2023tapir] with equal probability. When training with our dataset, we do not supervise the occlusion prediction for points identified as occluded, as we intentionally filter the dataset for high precision. We augment the filtered dataset using affine transformations, following the approach used in BootsTAP [doersch2024bootstap].
Optimization is performed using AdamW [loshchilov2017decoupled], with the learning rate of $3\times 10^{-4}$ and weight decay set to $1\times 10^{-3}$. The model is trained for 50K steps. We employ a cosine learning rate schedule with a 1000-step warmup, and apply gradient clipping with a maximum norm of 1.0. Training is conducted with a batch size of 1 per GPU, using 4 NVIDIA A6000 GPUs, and converges within 1 day. This is significantly more efficient compared to CoTracker3, which uses 32 A100 80GB GPUs for training, and BootsTAPIR, which uses 256 A100 GPUs. For each batch, we randomly sample 256 tracks and use a resolution of $256\times 256$.
Evaluation protocol. We use the TAP-Vid [doersch2022tap] benchmark to evaluate our model trained on Anthro-LD. We also evaluate our model on the RoboTAP dataset [vecerik2024robotap], a real-world robot dataset featuring videos of robotic manipulation tasks. The benchmark uses two evaluation modes based on how query points are selected: the First setup samples the query point from the first visible frame of each trajectory, whereas the Strided setup samples the query point every five timesteps along the trajectory. For evaluation, we use the following metrics, proposed in Doersch et al. [doersch2022tap]: Average Jaccard (AJ), position accuracy ($\delta_{avg}^{x}$), and Occlusion Accuracy (OA).
Trajectory metrics. To characterize the motion captured by our pseudo-labeled trajectories $\mathcal{X}_{p,j}$, we use simple metrics for measuring their complexity and diversity. Trajectory complexity is measured using the mean angular acceleration, which reflects how sharply and frequently a trajectory changes direction, computed over all contiguous visible segments. Diversity is assessed by evaluating how much individual trajectories deviate from a mean centered trajectory within each video, providing a measure of shape variability independent of absolute positions. Together, these metrics summarize how dynamic and varied the trajectories are across the dataset. Refer to Sec. D of the supplementary material for details.
4.2 Main results
Quantitative comparisons.
In Table 1, we present quantitative results on the TAP-Vid benchmark using the LocoTrack [cho2024local] base model trained with the dataset generated by our pipeline. Our primary comparisons are against BootsTAPIR [doersch2024bootstap] and CoTracker3 [karaev2024cotracker3] on the DAVIS [Perazzi2016] and Kinetics [kay2017kinetics], which consist of real-world videos. BootsTAPIR and CoTracker3 were trained on 15 million and 15,000 real videos respectively, whereas our model is trained on just 1,400 videos.
Despite this significant difference in training data scale, our method achieves notable performance gains. On the DAVIS First subset, our model surpasses BootsTAPIR by 3.3 percentage points on the $\delta^{x}_{avg}$, demonstrating the data efficiency of our approach. Additionally, it outperforms CoTracker3 by 2.0 percentage points on the same metric.
Table 2: Can training on human points generalize to non-human points? To answer this question, we compare the performance by grouping query points on humans and non-human regions separately in TAP-Vid-DAVIS dataset [doersch2022tap]. Our method shows greater improvement on non-human points.
Benchmark Method AJ $\delta^{x}_{avg}$ OA
DAVIS (Human Only) LocoTrack 50.7 42.4 57.8
Anthro-LocoTrack (Ours) 51.2 (+0.5) 43.3 (+0.9) 58.6 (+0.8)
DAVIS (Non-Human) LocoTrack 58.8 73.1 83.9
Anthro-LocoTrack (Ours) 60.8 (+2.0) 75.2 (+2.1) 85.3 (+1.4)
Table 3: Comparison to self-training method introduced in CoTracker3 using the same video. We compare our pipeline with self-training introduced in Karaev et al. [karaev2024cotracker3]. (I) CoTracker3 model trained on Kubric [greff2022kubric]. (II) fine-tunes the model on the Let’s Dance (LD) dataset using the self-training introduced in the paper. (III) trains CoTracker3 model using our pipeline. Our method shows greater performance boost than self-training introduced in CoTracker3.
Model Train Strategy Dataset DAVIS
AJ $\delta^{x}_{avg}$ OA
(I) CoTracker3 Supervised Kubric 63.8 76.3 90.2
(II) CoTracker3 Self-training LD [castro2018let] 64.2 (+0.4) 76.5 (+0.2) 89.6 (-0.6)
Qualitative comparisons. In Figure 3, we visually compare the results of CoTracker3 [karaev2024cotracker3], BootsTAPIR [doersch2024bootstap], and LocoTrack trained with our method. Our model demonstrates more robust tracking performance even in drastic deformation and severe occlusions. Refer to Sec. B of the supplementary material for additional visualization.
4.3 Ablation and analysis
Can training on human points generalize to non-human points?
In Table 4.2, we investigate this question by splitting the query points in the TAP-Vid-DAVIS dataset into human and non-human categories, and comparing the performance gain in each group relative to the baseline. We use Mask2Former [cheng2021mask2former] to identify points on humans. The results show that the performance improvements are significant for both human and non-human points, with an even greater boost observed for non-human points.
Comparison with self-training strategy of CoTracker3 on identical data. Table 4.2 compares our approach with the self-training strategy of CoTracker3 [karaev2024cotracker3] on identical data. We fine-tuned the Kubric-pretrained CoTracker3 baseline on the Let’s Dance dataset [castro2018let], the same data our pipeline employs, via two methods: the original self-training from CoTracker3, and fine-tuning with pseudo-labels from AnthroTAP. This allowed a direct comparison of both data utilization strategies on the same baseline and data.
Results in Table 4.2 show self-training by CoTracker3 yielded marginal baseline improvement on Let’s Dance. In contrast, pseudo-labels from AnthroTAP achieved significantly better performance, demonstrating that the quality of training signals from real-world data is important, not merely the use of such data.
Comparative analysis on trajectory complexity and diversity. In Table 4.3, we provide a comparative analysis of prevalent point tracking training datasets, focusing on trajectory complexity and diversity. We quantify trajectory complexity using the mean angular acceleration for each trajectory, thereby accounting for variations in direction. Trajectory diversity is assessed on a per-video basis by computing the mean standard deviation within each video. Anthro-LD, the Let’s Dance dataset [castro2018let] annotated using the AnthroTAP pipeline, shows the highest values for complexity and demonstrates greater diversity compared to other real-world datasets, such as DriveTrack [balasingam2024drivetrack].
Table 4: Comparison of track complexity and diversity with existing datasets. Our dataset exhibits significantly higher complexity and substantially greater diversity compared to another real dataset, DriveTrack.
Training Dataset Data Type Traj Metrics
Complexity Diversity
DriveTrack [balasingam2024drivetrack] Real 0.4396 0.0073
PointOdyssey [zheng2023pointodyssey] Synthetic 0.5222 0.1597
Kubric [greff2022kubric] Synthetic 0.1772 0.1165
Anthro-LD (Ours) Real 1.2492 0.1008
Figure 4: Qualitative comparison of trajectory complexity. Anthro-LD, a dataset generated from our pipeline, exhibits diverse and complex motions, unlike the rigid and simple trajectories found in DriveTrack [balasingam2024drivetrack].
Comparing real-world pseudo-GT with synthetic perfect GT. In Table 5, we investigate the relative importance of GT accuracy versus real-world training signals using BlinkVision [li2024blinkvision], a synthetic human mesh generator providing perfect GT with human motion. We train LocoTrack on three configurations: baseline, BlinkVision (perfect GT, synthetic), and AnthroTAP annotated dataset (pseudo-GT, real-world). Results show BlinkVision improves over baseline, but AnthroTAP annotated dataset achieves substantially larger gains despite pseudo-GT labels. Results show BlinkVision improves over baseline, while AnthroTAP annotated dataset achieves substantially larger gains despite pseudo-GT labels. Both datasets contain similar human motion patterns, differing primarily in GT accuracy and real-world visual characteristics.
Table 5: Comparison of models trained using point tracks sampled from synthetic human videos. The human points sampled from synthetic human videos do not provide a significant improvement over the baseline.
Method Datasets DAVIS
AJ $\delta_{avg}$ OA
LocoTrack ($\times256$) Kub64 63.0 75.3 87.2
LocoTrack + BlinkVision [li2024blinkvision] Kub64 + 1.4K 64.2 76.4 88.8
LocoTrack + Ours Kub64 + 1.4K 64.8 77.3 89.1
Impact of trajectory complexity and diversity on training. In Table 6, we investigate how trajectory complexity and diversity affect point tracking performance by creating partitioned datasets based on the metrics defined in Table 4.3. We train LocoTrack on three configurations: baseline, low-complexity trajectories only, and full dataset. Low-complexity trajectories are defined as those below the median angular acceleration value in our dataset. Results show that training on the full dataset, which contains high-complexity and high-diversity trajectories, achieves the largest performance gains. Training on low-complexity trajectories alone provides marginal improvements and even decreases performance on certain metrics compared to baseline. This indicates that trajectory complexity and diversity from human motion provide valuable training signals for point tracking models.
Table 6: Analysis of the impact of track diversity and complexity. We evaluate how training with low-diversity and low-complexity tracks affects model performance.
Method DAVIS
AJ $\delta_{avg}$ OA
LocoTrack 67.8 79.6 89.9
LocoTrack w/ Least complex tracks 66.0 79.9 86.9
LocoTrack w/ Least diverse tracks 66.6 79.8 87.9
Anthro-LocoTrack ($\times256$) 69.0 81.0 90.8
Ablation on optical flow based track filtering. In Table 4.3, we ablate the effect of optical flow-based filtering by comparing it with the baseline that uses trajectories projected from the human mesh and occlusion prediction via ray casting. While the baseline already achieves strong performance, applying trajectory rejection yields a further performance boost.
Visual comparison with DriveTrack [balasingam2024drivetrack]. In Figure 4, we qualitatively compare our dataset with DriveTrack [balasingam2024drivetrack], which is also a point-tracking dataset sourced from real videos. Observing the point traces, DriveTrack mostly exhibits monotonic motion, and the tracks tend to follow similar trajectories because they are concentrated on rigid objects undergoing largely uniform movement. In contrast, our dataset contains highly complex and diverse motion patterns.
Table 7: Ablation on optical flow based track rejection. Optical flow based rejection helps the model to learn more refined tracks.
Method DAVIS
AJ $\delta^{x}_{avg}$ OA
Baseline 64.4 76.9 88.6
+ Optical flow based rejection 64.8 77.3 89.1
5 Conclusion
In this paper, we presented a novel method for generating highly complex pseudo-labeled data for point tracking by leveraging the inherent complexities of human motion captured in videos. By fitting SMPL models to real-world human videos, accurately modeling occlusions via ray-casting, and refining trajectories using optical flow consistency, our approach significantly alleviates the bottleneck of manual annotation. A model trained on our dataset achieves state-of-the-art results with orders of magnitude less data than existing methods. Our method provides an effective path forward for robust point tracking in real-world scenarios.
Appendix
Appendix A Additional ablations and analyses
Table 8: Ablation on training video length.
Video Length DAVIS
AJ $\delta^{x}_{avg}$ OA
(I) $24$ 64.7 76.8 89.4
(II) $48$ 64.8 77.3 89.1
(III) $64$ 64.5 77.1 88.9
Ablation on training video length. In Table A, we investigate the influence of training video length on model performance. For this analysis, we fine-tuned the LocoTrack [cho2024local] base model using our annotated videos, experimenting with distinct clip lengths. The results demonstrate that utilizing video clips of 48 frames yields the optimal performance in terms of both AJ and $\delta^{x}_{avg}$.
Table 9: Ablation on frame dilation.
Frame DAVIS
Dilation AJ $\delta^{x}_{avg}$ OA
(I) 1 64.6 76.7 89.6
(II) 2 64.8 77.3 89.1
(III) 3 64.2 76.9 89.0
[…]
Table 10: Ablation on using different amount of extra data.
Method DAVIS
AJ $\delta^{x}_{avg}$ OA
LocoTrack ($384\times 512$) 64.8 77.4 86.2
LocoTrack + Ours (250 videos) 65.6 78.5 87.3
LocoTrack + Ours (500 videos) 65.7 78.7 87.2
LocoTrack + Ours (full) 65.9 78.9 87.3
Ablation on training data scale. In Table A, we examine how the amount of additional training data influences model performance. Starting from our full dataset containing 1.4K videos, we evaluate reduced subsets of 500 and 250 videos. The results show that even when using only a fraction of the full dataset, our additional videos consistently improve performance over the LocoTrack baseline across all metrics. This demonstrates that our dataset provides strong supervision signals, and meaningful gains can be achieved even with limited data.
Table 11: Ablation on proposed visibility prediction method.
Method DAVIS
AJ $\delta^{x}_{avg}$ OA
LocoTrack ($384\times 512$) 64.8 77.4 86.2
LocoTrack + Ours (w/o visibility prediction) 64.6 77.7 85.2
LocoTrack + Ours (full) 65.9 78.9 87.3
Ablation on visibility prediction. In Table A, we evaluate the impact of our proposed visibility prediction module. Removing the visibility signal results in a performance drop in both AJ and OA, indicating that visibility estimation provides useful supervision during training. The full pipeline with integrated visibility prediction achieves the best performance across all metrics.
Appendix B Additional visualization
We visualize videos annotated using our pipeline in Figure 7 and Figure 8. The visualizations highlight the complex motion present in the dataset.
Appendix C Filtering outlier trajectories with optical flow
@@
While HMR methods [black2023bedlam, dwivedi2024tokenhmr, goel2023humans] have become increasingly robust to challenges such as motion blur, rapid movements, and varying lighting conditions due to their strong human priors, they may still produce occasional errors or inconsistencies in mesh predictions under extreme conditions. Furthermore, our ray-casting-based visibility prediction is limited to occlusions caused by other modeled humans and does not account for occlusions from general scene objects. In addition, the parametric nature of the SMPL model makes it less effective at capturing the motion of highly deformable or loosely attached clothing and accessories.
To mitigate the impact of these potential inaccuracies on our pseudo-labels and thus improve the quality of training data, we introduce an optical flow-based filtering stage, illustrated in Figure 5. Optical flow [teed2020raft, huang2022flowformer, wang2024sea] is particularly well-suited for this task due to its established accuracy in estimating dense motion between adjacent frames. Over very short temporal windows, these flow-based predictions can provide strong local motion cues to validate or identify discrepancies in the HMR-derived trajectories, especially when dealing with complex non-rigid deformations or potential tracking drift [cho2024flowtrack, Neoral_2024_WACV].
@@
Identify confident optical flow. For each pair of consecutive frames $(I_{t},I_{t+1})$, we compute both the forward optical flow map from $I_{t}$ to $I_{t+1}$, denoted $\mathbf{F}_{t\to t+1}$, and the backward optical flow map $\mathbf{F}_{t+1\to t}$. We use $\mathbf{f}(\mathbf{F},\mathbf{x})$ to denote the process of sampling the flow vector from map $\mathbf{F}$ at sub-pixel location $\mathbf{x}$ using bilinear interpolation. For notational simplicity, we define:
$\bm{\xi}_{p,t,j}$ $=\mathbf{f}(\mathbf{F}_{t\to t+1},\mathbf{x}_{p,t,j}),\quad$ (1)
$\bm{\zeta}_{p,t,j}$ $=\mathbf{f}(\mathbf{F}_{t+1\to t},\mathbf{x}_{p,t,j}+\bm{\xi}_{p,t,j}).$ (2)
To assess the reliability of the optical flow estimation itself at a given point $\mathbf{x}_{p,t,j}$, we apply a forward-backward consistency check [meister2018unflow]. Let $c_{p,t,j}$ denote the binary indicator of optical flow reliability at point $\mathbf{x}_{p,t,j}$ in frame $I_{t}$. A pixel $\mathbf{x}_{p,t,j}$ is considered to have reliable flow ($c_{p,t,j}=1$) if the $L_{2}$ distance between the point warped to frame $t+1$ and then back to frame $t$ is below a predefined threshold $\delta_{\text{cons}}$:
$c_{p,t,j}=\mathbbm{1}\left[\|\bm{\xi}_{p,t,j}+\bm{\zeta}_{p,t,j}\|_{2}<\delta_{\text{cons}}\right].$ (3)
Find erroneous trajectories. For each point $\mathbf{x}_{p,t,j}$ within a pseudo-labeled trajectory $\mathcal{X}_{p,j}$ (where $t$ and $t+1$ are both in $\mathcal{T}_{p,j}$), we examine the predicted motion to the next frame, $\mathbf{x}_{p,t+1,j}$. The displacement vector derived from the HMR pseudo-label is $\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}=\mathbf{x}_{p,t+1,j}-\mathbf{x}_{p,t,j}$. We compare this to the optical flow displacement $\bm{\xi}_{p,t,j}$ (previously defined as the forward flow at location $\mathbf{x}_{p,t,j}$).
To compare these two displacement vectors, $\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}$ and $\bm{\xi}_{p,t,j}$, robustly, especially in cases of large motion, we normalize them. Let $L_{\text{short}}=\min(\|\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}\|_{2},\|\bm{\xi}_{p,t,j}\|_{2})+\epsilon_{\text{norm}}$, where $\epsilon_{\text{norm}}$ is a small positive constant added to prevent division by zero. The normalized vectors are:
$\Delta\widehat{\mathbf{x}}_{p,t,j}^{\text{HMR}}=\frac{\Delta\mathbf{x}_{p,t,j}^{\text{HMR}}}{L_{\text{short}}},\quad\widehat{\bm{\xi}}_{p,t,j}=\frac{\bm{\xi}_{p,t,j}}{L_{\text{short}}}.$ (4)
@@
$e_{p,t,j}=\mathbbm{1}\left[\|\Delta\widehat{\mathbf{x}}_{p,t,j}^{\text{HMR}}-\widehat{\bm{\xi}}_{p,t,j}\|_{2}>\tau_{\text{dist}}\right].$ (5)